OpenStudy (anonymous):

The length of the top of the table is 3m greater than the width. The area is 70m^2. Find the dimensions of the table.

OpenStudy (across):

\[l=3+w,\]\[lw=70.\]Do you know how to solve this system?

OpenStudy (anonymous):

No i do not sorry im terrible with math

OpenStudy (across):

\[(3+w)w=70,\]\[w^2+3w=70,\]\[w^2+3w+\frac{9}{4}=70+\frac{9}{4},\]\[\left(w+\frac{3}{2}\right)^2=\frac{289}{4},\]\[w+\frac{3}{2}=\pm\frac{17}{2},\]\[w=7,-10.\]

OpenStudy (anonymous):

Oh well thanks for showing me the steps and process :)

OpenStudy (across):

A table can't have a length of negative value, thus you have to pick 7.

OpenStudy (across):

no probs